OpenStudy (anonymous):

What is fraction 4/5 written as a percent

OpenStudy (anonymous):

A. 0.8% B. 8% C. 80% D. 800%

OpenStudy (anonymous):

\[4/5 = 8/10 = 80/100 = 80%\]

OpenStudy (anonymous):

Whoops, forgot the percentage sign. 80%